# Wake Vision: A Large-scale, Diverse Dataset and Benchmark Suite for TinyML Person Detection

### v1

@@ Appendix 0.F Label Generation Details
Appendix 0.G Standardized Evaluation
@@
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in Sec. 4.
Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier***Also known as alpha. of 0.25 [31]. For model-centric contributions we propose the following pre-processing pipeline:
@@
Appendix 0.H Dataset Download
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in Tab. 7.
@@ Appendix 0.H Dataset Download
Appendix 0.I Person Label Classes
@@
Appendix 0.J Wake Vision Dataset Size
@@
Image Level Label Train Dataset 2,880,214 2,880,214 2,176,551
Bounding Box Label Train Dataset 624,115 624,115 6,688,749
Validation Dataset 8,482 8,482 19,442
Test Dataset 25,641 25,641 58,023

### v2

@@ Appendix I Label Generation Details
Appendix J Standardized Evaluation
@@
To allow for both data-centric and model-centric improvements, we provide a standard model for data-centric contributions, and a standard pre-processing pipeline for model-centric contributions. Both types of contributions are expected to use accuracy as the primary metric for the overall test and validation set, and F1-score as the primary metric for the fine grained benchmarks introduced in section 4.
Therefore, for data-centric contributions we propose to use a Mobilenet v2 model with a width modifier [^1] Also known as alpha. of 0.25 [26]. For model-centric contributions we propose the following pre-processing pipeline:
@@
Appendix K Open Images Download
The full Open Images v7 dataset is not hosted by the dataset authors. Rather it is provided as a collection of flickr image Uniform Resource Locators (URLs) and their associated labels. As an unfortunate result of this, the dataset is not static over time as image owners can delete their images from the flickr platform. As a result we were only able to download a subset of the original Open Images v7 dataset as shown in table 6.
@@
Appendix L Person Label Classes
@@
Appendix M Wake Vision Dataset Size
@@
Wake Vision (Large) Training Dataset 2,880,214 2,880,214 2,176,551
Wake Vision (Quality) Training Dataset 624,115 624,115 6,688,749
Validation Dataset 9,291 9,291 17,824
Test Dataset 27,881 27,881 53,543
Appendix N Automatic Label Correction
To address the challenge of correcting Wake Vision validation and test labels we initially employed the Confident Learning technique [23] to intelligently identify potential label errors. We selected Confident Learning as prior work has demonstrated its capability to find label errors in large datasets [24]. Confident Learning flagged suspected mislabeled instances, which we inspected and corrected through a manual verification process.
Dataset Total Suggested Corrected Suggestion
Split Size Errors Errors Accuracy
Validation 18,582 632 81 12.82%
Test 51,282 1672 267 15.97%
As shown in Table 8, the confident learning process identified a large amount of possible label errors in Wake Vision’s Validation and Test sets. However, only between 12 and 16% of these possible label errors were legitimate errors. We corrected these identified label errors in the Wake Vision validation and test sets.
Given this low acceptance rate of label issues identified by Confident Learning, we concluded that we could not automate label cleaning to a point where no human in the loop is needed. This made this strategy too human-intensive to be applied to the much larger training sets.
To further correct label errors in the final Wake Vision Validation and Test sets we employed the Scale AI platform to crowd-source manual label corrections. These label corrections are described in section 3.2 and deprecate the automatic label corrections described in this section.